OpenStudy (anonymous):

solve the inequality: \left| x - 1 \right| + \left| 2- x \right| > 3 + x

OpenStudy (anonymous):

\[\left| x - 1 \right| + \left| 2- x \right| > 3 + x\]

OpenStudy (ash2326):

We have \[|x-1|+|2-x|>3+x\] There are two mods here so there are three conditions here Let's solve for each case seperately

OpenStudy (ash2326):

@kirmats are you here

OpenStudy (anonymous):

yupp...

OpenStudy (ash2326):

Case 1) x<1 Can you write the inequality for this condition

OpenStudy (ash2326):

@kirmats Are you working on it?

OpenStudy (anonymous):

1-x+ 2-x > 3 + x 3-2x>3+x -3x>0 x<0

OpenStudy (ash2326):

Good work:D So we started with the condition x< 1 and got x<0 so for this case what's the range of x?

OpenStudy (anonymous):

\[\left( -\infty, 0 \right)\]

OpenStudy (ash2326):

Great work:D Let's work on case 2 now Case 2) \( x= [1, 2)\)

OpenStudy (ash2326):

Could you solve for this case?

OpenStudy (anonymous):

x-1+2-x>3+x 1>3+x x<-2 this is where i got stuck..

OpenStudy (ash2326):

Ok, I'll explain We had the condition x<2 and x>=1 but we got x<-2 so in this case there is no common range, so no solution exist for this case range